# Co-Existence of Private 5G Network and Wireless Hospital Systems

### v1

@@ IV Methodology
This section presents the methodological framework for evaluating the coexistence of wireless devices and electromagnetic safety in hospital settings. The design prioritizes accuracy and reproducibility, conforming to established practices while mitigating challenges inherent to clinical environments. This approach facilitates the integration of technologies, such as private 5G without jeopardizing patient safety or disrupting incumbent medical systems. It aligns with IEC 60601-1-2 [iec_emc], ITU-T K.91 [itu_k91], and ICNIRP guidelines [ICNIRP2020], extending methodologies from prior RF studies in hospitals [rf_meas_hospital, coexistence_ieee2021].
@@ V-C Electromagnetic Safety and RF Exposure
The electromagnetic safety of the private 5G network in the operating room was thoroughly assessed. RF exposure levels in the 3.9–4.1 GHz band were found to be well below internationally recognized safety thresholds. The maximum received power of -16 dBm corresponds to a power density of approximately 0.05–0.1 mW, which is far below the limits set by ICNIRP and WHO for general public exposure [itu_emc, who_emc]. Even under worst-case assumptions (isotropic reception), exposure remains within safe limits, confirming the negligible biological risk posed by the in-hospital deployment of 5G. These results align with previous studies, indicating that properly engineered 5G systems do not pose significant health risks in hospital settings [WHO2020, ICNIRP2020].
@@
Furthermore, the assessment included both temporal and spatial variations in the operating room, ensuring that short-term peaks as well as cumulative exposure over the 24-hour measurement period remained within safe boundaries. This comprehensive evaluation is particularly important in hospital environments, where patients and staff may spend extended periods near active 5G transmitters. Fig. 5 illustrates that the peak received power remains comfortably within the green zone of RF exposure safety limits, significantly lower than the typical general public safety limit of 10 mW for frequencies above 400 MHz [ICNIRP2020]. The results provide strong evidence that the private 5G network operates safely under real-world conditions, ensuring both patient and staff safety during clinical operations.
@@ VI-A Spectrum Management and Practical Implications
From a practical standpoint, the results of this study provide valuable insights into spectrum management strategies for private 5G networks in hospitals. Adaptive sensing and thresholding techniques can be employed to dynamically adjust channel sensing based on real-time spectrum conditions. For example, the system can be adjusted to focus on stronger signals while filtering out background noise, allowing hospitals to minimize false detections and improve the reliability of spectrum monitoring, as suggested in previous studies. [al2016characterizing]. This approach ensures that real-time spectrum monitoring can be both efficient and reliable, even in a dynamic hospital environment with fluctuating RF conditions.
Additionally, the study reinforces the importance of guard bands and planned separation between vital 5G channels and known high-activity bands. For example, the 5G band in the hospital was carefully separated from the 2.4/5 GHz ISM bands, which are heavily used by Wi-Fi and Bluetooth devices. These efforts are consistent with regulatory guidelines that require careful frequency planning and interference mitigation, ensuring that 5G transmissions do not spill into adjacent medical telemetry or Wi-Fi channels [ITU2025]. Moreover, this study supports the recommendation to avoid using ISM bands for mission-sensitive 5G services, as these bands are heavily congested and unpredictable, especially in hospital settings where the number of wirelessly connected devices is growing rapidly.
@@ VI-C Regulatory and Policy Implications
The results of this study have broader regulatory and policy implications, particularly in relation to the adoption of dedicated spectrum for private networks in industries, as they demonstrate that private 5G networks can operate safely and effectively in hospital environments. This study aligns with ongoing international efforts to promote private network adoption in sectors such as hospitals, aviation, and manufacturing [itu_k91]. By demonstrating the feasibility of 5G coexistence, this work contributes to shaping future spectrum policies that support innovative, secure, and safe 5G network deployments across vital strategic infrastructures. The findings also underscore the potential for spectrum policy reforms that could facilitate the deployment of 5G networks in hospitals without disrupting existing systems. This supports the broader push for shared licensed spectrum or local spectrum regimes that allow private network rollouts in essential sectors.

### v2

@@ IV Methodology
This section presents the methodological framework for evaluating the coexistence of wireless devices and electromagnetic safety in hospital settings. The design prioritizes accuracy and reproducibility, conforming to established practices while mitigating challenges inherent to clinical environments. This approach facilitates the integration of technologies, such as private 5G without jeopardizing patient safety or disrupting incumbent medical systems. It aligns with IEC 60601-1-2 [23], ITU-T K.91 [27], and ICNIRP guidelines [24], extending methodologies from prior RF studies in hospitals [22, 12].
@@ V-C Electromagnetic Safety and RF Exposure
The electromagnetic safety of the private 5G network in the operating room was thoroughly assessed. RF exposure levels in the 3.9–4.1 GHz band were found to be well below internationally recognized safety thresholds. The maximum received power of -16 dBm corresponds to a power density of approximately 0.05–0.1 mW, which is far below the limits set by ICNIRP and WHO for general public exposure [26, 39]. Even under worst-case assumptions (isotropic reception), exposure remains within safe limits, confirming the negligible biological risk posed by the in-hospital deployment of 5G. These results align with previous studies, indicating that properly engineered 5G systems do not pose significant health risks in hospital settings [19, 24].
@@
Furthermore, the assessment included both temporal and spatial variations in the operating room, ensuring that short-term peaks as well as cumulative exposure over the 24-hour measurement period remained within safe boundaries. This comprehensive evaluation is particularly important in hospital environments, where patients and staff may spend extended periods near active 5G transmitters. Fig. 5 illustrates that the peak received power remains comfortably within the green zone of RF exposure safety limits, significantly lower than the typical general public safety limit of 10 mW for frequencies above 400 MHz [24]. The results provide strong evidence that the private 5G network operates safely under real-world conditions, ensuring both patient and staff safety during clinical operations.
@@ VI-A Spectrum Management and Practical Implications
From a practical standpoint, the results of this study provide valuable insights into spectrum management strategies for private 5G networks in hospitals. Adaptive sensing and thresholding techniques can be employed to dynamically adjust channel sensing based on real-time spectrum conditions. For example, the system can be adjusted to focus on stronger signals while filtering out background noise, allowing hospitals to minimize false detections and improve the reliability of spectrum monitoring, as suggested in previous studies. [6]. This approach ensures that real-time spectrum monitoring can be both efficient and reliable, even in a dynamic hospital environment with fluctuating RF conditions.
Additionally, the study reinforces the importance of guard bands and planned separation between vital 5G channels and known high-activity bands. For example, the 5G band in the hospital was carefully separated from the 2.4/5 GHz ISM bands, which are heavily used by Wi-Fi and Bluetooth devices. These efforts are consistent with regulatory guidelines that require careful frequency planning and interference mitigation, ensuring that 5G transmissions do not spill into adjacent medical telemetry or Wi-Fi channels [25]. Moreover, this study supports the recommendation to avoid using ISM bands for mission-sensitive 5G services, as these bands are heavily congested and unpredictable, especially in hospital settings where the number of wirelessly connected devices is growing rapidly.
@@ VI-C Regulatory and Policy Implications
The results of this study have broader regulatory and policy implications, particularly in relation to the adoption of dedicated spectrum for private networks in industries, as they demonstrate that private 5G networks can operate safely and effectively in hospital environments. This study aligns with ongoing international efforts to promote private network adoption in sectors such as hospitals, aviation, and manufacturing [27]. By demonstrating the feasibility of 5G coexistence, this work contributes to shaping future spectrum policies that support innovative, secure, and safe 5G network deployments across vital strategic infrastructures. The findings also underscore the potential for spectrum policy reforms that could facilitate the deployment of 5G networks in hospitals without disrupting existing systems. This supports the broader push for shared licensed spectrum or local spectrum regimes that allow private network rollouts in essential sectors.